Restart 20150913 UT

OSA: Geno Bechetti ZZ ISp: Elliott Shield

Manager: Kevin Newton

ISAs: John Hill, Olga Kuhn, B. Rothberg (Tucson)

Overall Start-up Summary

MODS1 Summary from Olga

MODS1 is ready for science. Last night we obtained various data sets:
  • Spectra of the spectrophotometric standard, G191B2B, in all 6 spectroscopic modes (1 spec per mode since it was cloudy).
  • 6 dice5 sequences on the astrometric StoneO field.
  • Data to measure the rotator zeropoint offset (but we did not update this because it would require rotating the transform).
  • Long-slit (both for the 5" slit and narrower slits) and MOS alignments.
  • Data to measure the instrument rotator center.

I watched the data as they came in and did some quick analysis of some, and I see nothing unusual. Keep an eye on the long-slit alignments, as the offset seems to land the target about a pixel to the right of slit center. I will have to look at the 5" slit X-centers - objects seemed also to be a tad to the right of center here too, probably for a different reason, but this is not unlike what I have noticed in the past.

Prior to sky time, I updated the IMCS TTF zeropoints (minor change but we do this each year) and took quite a few calibrations. The offset needed to the rotator zeropoint was measured to be -0.2 deg, based on the slitscan sequence, and this is consistent with the MOS alignments which indicated a rotational offset of ~ -0.15 deg needed. No action was taken, however, since the transform would need to be updated to be consistent with an updated rotator zeropoint.

  • On the guiding/WFSing: John and Doug collected and analyzed the field aberration data, but have not yet verified the results. I would not expect that this would make a big difference, however. So I believe that MODS1 is ready for science. There are some improvements that could have been made with more sky time (e.g. the rotator zeropoint) and investigation (slit centering), but it is behaving as it did earlier.

LBC Summary from Barry

LBC-wise, I think it should be OK. We still have not had a chance to take skyflats, but we have run through all filters and dithering with and without guiding. We noticed an issue in collimation, but it turns out the hub collimation table that was updated for the summer position was not recovered after the SANS crash, so I just updated the files again (this explains some oddities in collimation we noticed the last few nights).

However, we have one issue that popped up last night that remains. We cannot dither LBC-Red while MODS1 is in the SX beam and we are in pseudo-monocular mode. Staring works fine though. This may be related to the latest TCS update.

DOFPIAlow seems to work OK. I would recommend the users actually use this now, but John Hill and Olga may want to chime in on that.

There are still some second tier things to be done with the LBCs, including checking ZPs (need photometric nights for that).

Night Summary

John and Geno aligned the rigid secondary using X and Y lateral displacements. John and Doug took an excellent set of field aberration measurements, but ran into some problems with the analysis and will defer that to later (Sunday morning). Barry and John ran LBCR in pseudo-monocular mode and tested dithering. There was a problem with the rotator trajectory after a dither (IT 5680). They used both the LBC current version 2.0.3 and the new version 2.1. I did basic checkout on MODS1:
  • Wide-Slit, Long-Slit and several MOS alignments using 2 MOS masks on StoneO and StoneA fields.
    • Long-slit acquisitions seemed to end up with stars just a bit, ~1pix, to the right. This is not new, as we've noticed this in the past semester(s).
    • Wide-slit alignments also seemed to have the star a bit to the right, probably for a different reason.
    • 2-3 MOS alignments for each mask give similar rotational and x,y offsets.
  • Took data to measure rotator center. (finally determined the imcslock should not be run between rotator angles...so got end-of-night set that looks better).
  • Took 'slitscan' data to measure offset of rotator zeropoint. The result, consistent the the rotational offset needed to align the MOS masks, about -0.2 deg.
  • Ran several dice5's on StoneO, with 10" and 1" offsets.
    • Used new LBTtools.Sandbox task, modsastrom, to put WCS on a few images. Pixel scales are 0.120"/pix and 0.123"/pix as expected.
    • Compare expected with actual offsets following method in MODS1 AGw Comm report.
  • Ran through all modes (1 image per mode) on a standard star (G191B2B). Clouds. Looks OK.

I think MODS1 is working as usual - nothing unusual to report. Will try to look harder at the data, but did watch it coming in and did some quick analysis (as noted above), and all looks OK.

Details

UT 03:16 - Opening

UT 03:22 - actually open, send ACQUIRE preset

LDG Field Aberrations for Rigid Secondary

UT 03:xx - ACQUIRE preset to Vega

UT 03:xx - rigid secondary aligned roughly by X, Y translation to recover pointing and coma.

UT 04:01 - AGW field collection started - called Doug to fix a spurious IDL prompt.

UT 04:02 - Olga carefully preparing MODS tests

UT 04:07 - BS9177 - for field collection resumes

Doug changed the pupil center during image wfsc 0071. Guider jump during image wfsc 0097.

Clouds rolling in after GS12 - see image wfsc 0101.

GStar wfsc #
2015913
BS9177_field
wfsc
GCS
0 59-60
11 61-63
0 68-69
16 70-72
0 73
13 74-76
0 77-78
17 79-81
0 82
21 83-85
0 86
6 87-89
0 90
3 91-93
0 94
4 95-97
0 98
12 99-101
0 102

Field Aberration Results

Field analysis results: Move secondary Tip by -12" and add 12" to pointing model IE

 => Field Data Results
             Measured bi-nodal astigmatism
  
                  ldg
              From 45 Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip = -21.1", Delta IE =  21.1"
         Tilt =  -4.1", Delta CA =   4.1"
  
  
                  ldg
              From 0 Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =  -8.0", Delta IE =   8.0"
         Tilt =   3.7", Delta CA =  -3.7"
 

Doug's -12 arcsec above make the SX Pointing term below.

SX Global X=-1.500 Y=-5.600 Z=-3.400 RX=0 RY=0
SX Active X=+0.794 Y=+2.830 Z=+0.084 RX=-294.9 RY=+82.7
SX Pointing X=0.000 Y=+0.613 Z=0.000 RX=-117.8 RY=0.0


SX Total X=-0.706 Y=-2.157 Z=-3.316 RX=-412.7 RY=+82.7

This will be added to SXSMLUCIFERCollimation.20140322.dat to create SXSMLUCIFERCollimation.20150914rigid.dat which was propagated to all the SX focal stations on 20150914 UT.

Added 12 IE arcsec to SX_Gregorian.20131218.ptmod to make SX_Gregorianrigid.20150914.ptmod. Added 12 IE arcsec to SX_PFU.20150524.ptmod to make SX_PFUrigid.20150914.ptmod.

MODS1 Checkout and Calibration

UT 04:10 - started MODS1 slit alignment check while Doug and John working on IDL field collect script. First slit and field images had guide probe on-axis. On field image, stars looked round. Cntl-C'd the script as I homed the guide probe. Repeated script, which should have done 'instconfig dual imaging', but next 60-sec slit images, blue and red, (0003) are clearly stepped. Not sure what was going on. When taking a series of slit images, I noticed, for blue where exptime was just above threshold for imcslock, that the series of slit images were shifted with respect to one another, not lined up on the same columns as they were for red, where exptime was too low to trigger closed-loop IMCS operation. IMCS thinks it's locked, but some hard-to-explain shifts. It might have been that Doug and John were moving the rotator while I was taking images 3 on.

UT 05:23 - acqMODS mods.36.StoneO.acq
  • R=16.1 guide star was too faint for these conditions. Could not wfs. After 30-sec mask image (6, boxes look sharp), control-c'd script and resent preset with R=13.09 guide star instead (hadn't used it earlier because it will cause some vignetting of boxes in lower left). Much better.
  • mask = 7, field = 8
    • modsAlign mods.36.StoneO.mms mods1r.20150913.0007.fits mods1r.20150913.0008.fits, used 6 stars distributed about field for alignment:
1563.46 1557.14    471.
1595.07 2296.70    481.
2366.23 2650.67    469.
2511.61 1474.97    466.
1601.39  299.27    471.
 160.20 1753.09    458.

MODS Offset Command:
   offsetpointing -0.1549 -3.805 -0.321 detxy rel
  • confirm = 9, looks pretty good. Stars are in the boxes.

UT 05:40 - repeating acqMODS mods.36.StoneO.acq
1576.10 1549.77    455.
1818.41 2060.71    468.
 643.76 1391.74    461.
2503.18 1465.49    423.
 538.41  791.25    463.
One of these was the double star.

MODS Offset Command:
   offsetpointing -0.0463 -3.714 -0.473 detxy rel
  • confirm = 12. Stars are in boxes, but may not be as good as 9.

UT 05:53 - dice5 pattern on StoneO field. First 5 10" dithers, then 5 1" dithers.
  • The dice scripts had OBSMODE Imaging (this was not the problem).

UT 06:00 - execMODS mods.36.StoneO.obs (dual grating spectrum of mask).
  • while I figure out what is going on with the imaging dice script... OK, I'd copied the dice script from MODS2 Commissioning and forgot to change execMODS2 to execMODS (this was the test I did not prepare too carefully...).
  • Abort ongoing script while first of 3 images is reading out and during the offset 10 10 abs command. First images are b 6 and r 13.

UT 06:05 - runDice5 (on MODS1!) on StoneO. This will repeat a 10" dice-5 three times, then a 1" dice-5 three times. (Remove mask and restart).
  • modsastrom on first images -> pixel scales we expect.
  • on b 12 and r 19, wfs fwhm 1.34". MODS images also look elongated.
  • compared first and 2nd offset 0 0 images and they overlap well.
  • Losing guide signal now: was R=14, now R=16-17. Note that guide star should have R=13.09, so we may need to take into account a a 1 mag offset for LBTplot. This is for the MODS1 guider, clear filter.
  • Note that although offset command is for RADec absolute offsets, offset 10.0 10.0 abs, GUI reports CoordSys RADec and MoveType Relative. I believe this is normal behavior

UT 07:05 - Barry and John testing lbc version 2.1 on LBC-Red.

UT 07:07 - runDice5 more to do 2 more dice10 and dice1 (execMODS -e).
  • b 40 and r 48 - guiding is lost.

UT 07:12 - Closing for clouds.

UT 08:44 - Open again.

UT 08:45 - mask alignment on 2nd Stone mask (mods.37.StoneA.mms). Used the boxes at:
1307.84 1326.05    358.
1772.06 1926.79    373.
1626.42 2309.09    380.
 279.29  470.44    365.
2236.27  634.28    369.
2336.39  251.98    370.
2891.63 1316.94    364.

MODS Offset Command:
   offsetpointing -0.1511 -3.837 -1.063 detxy rel

All objects are in the slits/boxes.

acqMODS -i and acqMODS -a to repeat sequence and do alignment one more time while setting up slitscan script. This time I did not repeat the preset, so the small offsets are understandable. But I don't understand the small rotational offset on the 2nd acq of the StoneO mask.

MODS Offset Command:
   offsetpointing -0.0478  0.365  0.065 detxy rel

confirm = 54 (mount jumped at this time).

UT 08:57 - slitscan on StoneA field --- aborted because clouds too thick and the guide star was not found.

UT 09:11 - acqMODS g191b2b.acq

MODS Offset Command:
   offsetxy -3.210  6.171 rel

confirm = 56. Star centered at 500.2, 587.6, so it was not quite at the sweet spot of 501.6, 590.

UT 09:20 - execMODS g191b2b_checkout.obs
  • 1 observation at each configuration. Tagged as (cloudy).
>> instconfig blue grating
** ERROR: Command 'instconfig blue grating' timed out after 60 seconds
** Abort, Retry, or Ignore? > r
** Retrying command 'instconfig blue grating'
>> instconfig blue grating
** ERROR: instconfig instrument configuration already in progress - command disallowed
** Abort, Retry, or Ignore? > 

I could not find the source of error - after a few minutes of searching, I tried to "r"etry again, and the instconfig blue grating command completed.

UT 09:50 - acqMODS g191b2b again, because in first acquisition, after the offset, the star was 1-2 pixels off. Guide star signal is very low and just disappeared.

UT 09:52 - Barry restored the version 2.03 of the LBC software.

UT 10:06 - Clouds have thinned. acqMODS -a g191b2b to retake the acquisition image.
MODS Offset Command:
   offsetxy -3.086  5.930 rel

But we had an offset error.
Send the offset to the Telescope (Y|N)?: Y
   Sending offset...
*** ERROR: offsetxy IIF error - OffsetPointing result status: Error, PCS OffsetTelescopeDETXY SYNCHRONIZED OffsetTelescope aborted on side left., PCS OffsetTelescopeDETXY failed on side left.
   Offset command returned error - /home/MODSeng/bin/modsAlign aborting

John thinks that LBC set a synchronized RADEC offset and MODS sent a synchronized detxy offset. MODS + LBCR pseudo-mono, but IIF did not have "ignore" sync offset button pressed. Barry checked this button, and I sent the offset again via the command window of the MODS GUI. Worked now.

Star at 502.6, 589.5 near right edge of slit.

UT 10:20 acqMODS g191b2b.acq.

Barry and John have been testing LBC (pseudo-mono) with LBC dithering. The rotator trajectory is not correct, and therefore guiding is lost, after the dither. This is current LBC version, v2.0.3. It's a TCS problem they think.

MODS Offset Command:
   offsetxy -3.184  6.177 rel
Send the offset to the Telescope (Y|N)?: 
   OK, not sending the offset...

confirm=65. I had just hit return and not sent the offset. Sent the offset through the command window of the GUI. confirm=66. 502,590. star is a bit to the right.

UT 10:28 Run slitscan on this field. Losing guide star again.

execMODS -f Instrument slitscang191.img to configure instrument and take slit image only.

execMODS slitscan2g191.img to send a track preset and to offset up/down along the slit (this slewed from g191b2b to StoneA).

UT 10:42 execMODS slitscan1a.img (StoneA).
  • slitimage 68
  • images 69-79
Line of star centroids is -0.663+/-0.007 degrees CCW from the Y axis
Slit is                   -0.411+/-0.000 degrees CCW from the Y axis 
-0.252+/-0.007 degrees should be added to the current rotator zeropoint 

UT 11:09 Try acquisition on StoneA again MOS mask again.
MODS Offset Command:
   offsetpointing -0.1171 -3.513 -1.763 detxy rel

confirm = 82.

acqMODS mods.37.StoneA.acq again to see whether the rotational offset will be close to zero or closer to previous value (does the preset to the same PA remove it)? No, it seems not, as expected. Earlier ~0 rot offset was anomaly.
MODS Offset Command:
   offsetpointing -0.1850 -3.488 -1.211 detxy rel
confirm = 85

UT 10:26 Long slit acquisition on G191B2B. 1" slit. execMODS g191b2b_1asec.acq. Guide star is in and out of the clouds, but WFS Err down to 400 nm before slit image taken. Measured slit center at Y=620 (to avoid the bad columns the spectrum goes through when centered at Y=590).
MODS Offset Command:
   offsetxy -3.129  8.996 rel
confirm = 88 star is a bit to the right of center.

UT 10:38: Rotator Center. Since we are already collimated on this target, send only the set of track presets at PA=0,180,90,270.
  • execMODS g191b2b_trk0.img (b 59 and r 89) and trk180 (b60 and r 90) (note, I really should not run the imcslock each time).
    • RED: (1562,1565) and BLUE: (1572,1553)
###===red=0 180=====================
Find rotation center with geomap...
   Xc=1561.97 Yc=1565.08 from geomap.
Find rotation center with coordinate averages...
   Xc=1562.19 Yc=1565.09 from averages.
###======blue 0 180 ========================
Find rotation center with geomap...
   Xc=1572.75 Yc=1552.93 from geomap.
Find rotation center with coordinate averages...
   Xc=1571.94 Yc=1553.39 from averages.
  • execMODS g191b2b_trk90.img and 270. (b 61,62 and r 91,92)
###=================red=90 270======
Find rotation center with geomap...
   Xc=1562.19 Yc=1565.61 from geomap.
Find rotation center with coordinate averages...
   Xc=1562.88 Yc=1565.16 from average
###=================blue 90 270==========
Find rotation center with geomap...
   Xc=1563.68 Yc=1560.15 from geomap.
Find rotation center with coordinate averages...
   Xc=1563.56 Yc=1559.97 from averages.

UT 12:00 Long slit acq on G191b2b again.
MODS Offset Command:
   offsetxy -2.963  8.759 rel
confirm = 95. Still offset to the right a bit. We have seen this before. Could the slit center be mesur

UT 12:08 Rotator center without imcs lock for each script.
  • slitmask imaging done manually
  • instconfig dual imaging done manually
  • imcslock done manually
  • execMODS g191_trk0.img (b 63 and r 97), g191_trk180.img (b 64 and r 98)
    • red: 1537,1581 and blue: 1557, 1568.
  • execMODS g191_trk90.img (b 65 and r 99), g191_trk270.img (b 66 and r 100)
    • red: 1531, 1575 and blue 1553, 1570.

UT 12:25 Closing. Put mods1 to sleep.

-- %USERSIG{BarryRothberg - 2015-09-13}%
Topic revision: r10 - 14 Sep 2015, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback